# Cross-Service Threat Intelligence in LLM Services using Privacy-Preserving Fingerprints

### v1

@@ 4.2 Attack Defenses
Yuan et al. (2024) propose RigorLLM framework that enhances content moderation by (a) enriching training data with diverse adversarial examples using constrained energy-based generation, (b) appending an optimized “safe suffix” to incoming prompts to mitigate adversarial modifications, and (c) fusing predictions from a robust KNN classifier and a fine-tuned LLM to reliably detect harmful content.  Hines et al. (2024) build on the work of  Yi et al. (2025) [^1] We cite the peer-reviewed conference version. An earlier preprint is available on arXiv. to address indirect prompt injection attacks by proposing a defense strategy based on spotlighting. This approach comprises a set of prompt-engineering techniques that transparently indicate the provenance of each input segment to assist LLM while generating response. The methods operate exclusively at inference time and do not require access to the model’s internals, making them readily applicable to API-based models.  Debenedetti et al. (2025) propose CaMeL, a defense mechanism that explicitly detects and prevents prompt injection attacks by designing a secure execution layer around LLM operations. The work sits at the intersection of generative AI and software security, bringing established principles from the latter to protect emerging systems. Abdelnabi et al. (2025a) proposes TaskTracker, white-box defense, that centers on tracking the model’s internal activations before and after processing external text. By calculating the change (or “delta”) between these activations, the system can flag when the model’s task representation has been altered by injected instructions. This drift is used as a proxy to signal potential prompt injections.  Liu et al. (2025) introduces DataSentinel that detects prompt injection attacks by fine-tuning a language model with a game-theoretic mini-max optimization. It uses a detection instruction embedding a secret key to verify the integrity of incoming data. Beurer-Kellner et al. (2025) propose a six architectural design patterns (action-selector, plan-then-execute, map-reduce, dual LLM, code-then-execute, and context-minimization) that secure LLM agents against prompt injection attacks. These patterns restrict an agent’s ability to process untrusted input with some trade-offs in agent flexibility.
Overall, most existing defenses are probabilistic and can be bypassed (Debenedetti et al., 2025; Costa et al., 2025). Retrospective systems are needed to identify attackers, yet cross-service threat intelligence for large-scale prompt injection attacks correlation remains unexplored.
4.3 Discussion
@@
Our study has limitations similar to those in any research. Synthetic paraphrasing and word substitution may yield varying results with different LLMs and prompts. The effectiveness of the BinaryShield’s PII redaction module depends on the accuracy of detection methods. Residual PII may remain in prompts. Our results rely on specific embedding models, ModernBert (open-source) (Warner et al., 2025; Gill et al., 2025) and OpenAI text-embedding-3-large (proprietary). Thus, performance may vary with other models. We determined optimal Hamming distance thresholds through search on our datasets. However, these thresholds may not generalize to other attack distributions or organizational contexts. All experiments contain English prompts. Future work should explore cross-lingual attack detection and domain-specific technical prompts. We depend on computational non invertibility and differential privacy guarantees. The evaluation does not assess resistance to reconstruction attacks, which represent an orthogonal concern to current focus of the paper.
@@
Noise can be added to the fingerprint through various differential privacy mechanisms (Warner, 1965; Wang et al., 2017; Bhowmick et al., 2018; Erlingsson et al., 2014; Feyisetan and Kasiviswanathan, 2021; Vadhan and Wang, 2021; Carvalho et al., 2021; Bollegala et al., 2023; Xie et al., 2024; Meehan et al., 2022; Asi et al., 2022; Du et al., 2023; Bittau et al., 2017; Fanti et al., 2016; Hsu et al., 2014; Near and Abuah, 2021; Fioretto et al., 2024; Dwork et al., 2014; Balle and Wang, 2018; Ding et al., 2017; Hu et al., 2024). BinaryShield utilizes foundational work of Randomized Response mechanism (Warner, 1965). Assessing the impact of other DP mechanisms and global differential privacy on the overall privacy guarantees is an important direction for future research. We did not formulate the privacy utility trade off as an optimization problem. Readers may consult existing literature on this topic (Near and Abuah, 2021; Wang et al., 2017; Dwork et al., 2014; Fioretto et al., 2024). However, in our experiments, we empirically evaluated conditions ranging from extreme privacy, where accuracy approaches zero, to moderate and weak privacy. Each organization has specific privacy policies, and we do not recommend a particular differential privacy budget. Organizations should work with their privacy teams to select an appropriate budget based on their guidelines. Note we do not claim any novelty in the individual techniques used in BinaryShield. Our primary contribution is the proposal of the first threat intelligence system for prompt injection threat correlation. We integrate insights from data anonymization, natural language processing, quantization, and differential privacy to design a practical and deployable threat intelligence system. This work constitutes an initial step in the field, and we hope it will inspire further research in cross-service threat intelligence for LLM misuse.
5 Conclusion
The proliferation of LLMs across enterprise services has created an unprecedented security challenge: organizations operate multiple LLM services with billions of daily queries, yet regulatory boundaries prevent these services from sharing critical threat intelligence about prompt injection attacks. This fragmentation leaves each service vulnerable to attacks that may have already been detected elsewhere within the same organization, creating a fundamental gap in collective defense capabilities.
@@ 5 Conclusion
BinaryShield addresses this critical vulnerability by introducing the first privacy-preserving fingerprinting mechanism specifically designed for prompt injection threat intelligence. Through a carefully orchestrated pipeline combining PII redaction, semantic embedding, binary quantization, and randomized response, BinaryShield generates non-invertible fingerprints that can be safely shared across compliance boundaries while preserving the semantic characteristics necessary for effective threat correlation. When compared to SimHash, the privacy-preserving baseline, BinaryShield demonstrates superior robustness. The operational efficiency gains are equally compelling.
@@
The implications of this work extend beyond immediate operational benefits. As LLM-based systems become critical infrastructure across finance, healthcare, and government services, the attack surface for prompt injection continues to expand exponentially. The emergence of autonomous agents and MCP servers further amplifies these risks, potentially enabling attacks that cascade from text manipulation to arbitrary code execution and system compromise. BinaryShield establishes the foundational infrastructure for collaborative defense against these evolving threats, creating a pathway toward industry-wide threat intelligence feeds analogous to existing malware signature ecosystems.

### v2

@@ IV Related Work and Discussion
Attack Defenses. Recent work proposes a variety of defense strategies against prompt injection attacks. These approaches differ in their methods and operating settings, ranging from inference-time countermeasures to white-box techniques and architectural safeguards. RigorLLM [56] framework enhances content moderation. Hines et al. [25] build on the work of  [55] to address indirect prompt injection attacks by proposing a defense strategy based on spotlighting. CaMeL [13] is a defense mechanism that explicitly detects and prevents prompt injection attacks by designing a secure execution layer around LLM operations. The work sits at the intersection of generative AI and software security, bringing established principles from the latter to protect emerging systems. SafeEar [32] proposes a method for detecting deepfake audio without accessing the actual speech content. TaskTracker [1] is a white-box defense, that centers on tracking the model’s internal activations before and after processing external text. DataSentinel [34] detects prompt injection attacks by fine-tuning a language model with a game-theoretic mini-max optimization. It uses a detection instruction embedding a secret key to verify the integrity of incoming data. Beurer-Kellner et al. [5] propose a six architectural design patterns (action-selector, plan-then-execute, map-reduce, dual LLM, code-then-execute, and context-minimization) that secure LLM agents against attacks. Overall, most existing defenses are probabilistic and can be bypassed [13, 12]. Retrospective systems are needed to identify attackers, yet cross-service threat intelligence for large-scale prompt injection attacks correlation remains unexplored. To address this gap, techniques like SimHash can be leveraged. However, SimHash captures only syntactic similarity and fails to capture semantic similarity (Figures 56, and 7). BinaryShield is the first to focus on this gap, proposing a concrete system for cross-service threat intelligence and paving the way for future research in this domain.
IV-A Discussion
@@
Our study has limitations similar to those in any research. Synthetic paraphrasing and word substitution may yield varying results with different LLMs and prompts. The effectiveness of the BinaryShield’s PII redaction module depends on the accuracy of detection methods. Residual PII may remain in prompts. Our results rely on specific embedding models, ModernBert (open-source) [50, 23] and OpenAI text-embedding-3-large (proprietary). Thus, performance may vary with other models. We determined optimal Hamming distance thresholds through search on our datasets. However, these thresholds may not generalize to other attack distributions or organizational contexts. All experiments contain English prompts. Future work should explore cross-lingual attack detection and domain-specific technical prompts. We depend on computational non-invertibility and differential privacy guarantees. The evaluation does not assess resistance to reconstruction attacks, which represent an orthogonal concern to current focus of the paper.
@@
Differential Privacy. Noise can be added to the fingerprint through various differential privacy mechanisms [51, 49, 6, 19, 21, 47, 10, 8, 53, 38, 3, 17, 7, 20, 26, 40, 22, 18, 4, 16, 27]. BinaryShield utilizes foundational work of Randomized Response mechanism [51]. Assessing the impact of other DP mechanisms and global differential privacy on the overall privacy guarantees is an important direction for future research. We did not formulate the privacy utility trade off as an optimization problem. Readers may consult existing literature on this topic [40, 49, 18, 22]. However, in our experiments, we empirically evaluated conditions ranging from extreme privacy, where accuracy approaches zero, to moderate and weak privacy. Each organization has specific privacy policies, and we do not recommend a particular differential privacy budget. Organizations should work with their privacy teams to select an appropriate budget based on their guidelines.
V Conclusion
Regulatory boundaries prevent an enterprise’s multiple LLM services from sharing prompt injection threat intelligence. This isolation means an attack blocked by one service can still succeed against another, as each service has its own defense mechanism. This creates a fundamental gap in the organization’s collective defense capabilities.
@@ V Conclusion
BinaryShield addresses this critical gap by introducing the first privacy-preserving fingerprinting mechanism specifically designed for prompt injection threat intelligence. Through a carefully orchestrated pipeline combining PII redaction, semantic embedding, binary quantization, and randomized response, BinaryShield generates privacy-preserving fingerprints that can be safely shared across compliance boundaries while preserving the semantic characteristics necessary for effective threat correlation.
@@
The implications of this work extend beyond immediate operational benefits. As LLM-based systems become critical infrastructure across finance, healthcare, and government services, the attack surface for prompt injection continues to expand exponentially. The emergence of autonomous agents and MCP servers further amplifies these risks, potentially enabling attacks that cascade from text manipulation to arbitrary code execution and system compromise. BinaryShield establishes the foundational system for collaborative defense against these evolving threats, creating a pathway toward industry-wide threat intelligence feeds analogous to existing malware signature ecosystems.
VI LLM Usage Considerations
We utilized GPT-4o to systematically generate the attack variants (word-flipped and paraphrased) for our evaluation. This approach was necessary because no public dataset exists for cross-service threat intelligence analysis and enterprise data access is restricted. To address this gap, we used GPT-4o to create a partially synthetic dataset from original prompt injection attacks. The generation was a one-time, offline process using efficient prompts (see Appendix LABEL:prompt:words-flipped and LABEL:prompt:paraphrase) to minimize the computational footprint. While using a closed-source model affects exact reproducibility, our generation process is transparent, and the resulting dataset characteristics are replicable with other powerful LLMs. We also briefly discuss this limitation in Section IV-A.
We also used LLMs to correct grammar mistakes, similar to a feature like Grammarly. The whole content in this study is written entirely by the authors, who take full responsibility for the paper.
Acknowledgment
We sincerely thank the SaTML reviewers for their valuable feedback. We are also deeply grateful to Yonatan Zunger, Angela Argentati, Dharmin Shah, Sukirna Roy, Habiba Mohamed, Anandan Sundar, Leah Zulas, and the entire Microsoft AI Safety and Security organization for their support.